solve the system by addition
5x-3y=13 4x-3y=11

Answer by askmemath(368)

Instead of adding try Subtracting
5x-3y = 13
-(4x-3y = 11)
5x-3y = 13
-4x +3y = -11
Adding
X = 2
Substituting X back in any equation
Y = -1